# Fourier Learning Machines: Nonharmonic Fourier-Based Neural Networks for Scientific Machine Learning

### v1

@@ 3.2 Cosine Phase-Shifted Basis for $m$-dimensional Input
In a similar fashion as in the two–dimensional case, $H_{\boldsymbol{n}}$ represents the linear combination of $l=2^{m-1}$ phase–shifted cosinescosines, which is equivalent to a linear combination of $2^{m}$ separable Fourier basis in (8) and can be written as
$H_{\boldsymbol{n}}=\sum_{i=1}^{l}A_{i}^{(\boldsymbol{n})}\cos\left((\boldsymbol{e}^{(i)}\odot\boldsymbol{n})\cdot\boldsymbol{x}-\phi_{i}^{(\boldsymbol{n})}\right),$
where $\boldsymbol{e}^{(i)}$ is the $i^{th}$ row of the $m$-Lexi Sign Matrix $S$ and $\boldsymbol{e}^{(i)}\odot\boldsymbol{n}$ represents its Hadamard (element wise) product with the frequency vector $\boldsymbol{n}$. The $m$-dimensional Fourier series can be expressed in a single compact summation as
$\widehat{f}_{NH}(\boldsymbol{x})=\sum_{\boldsymbol{n}\in\mathcal{N}}H_{\boldsymbol{n}}.$ (13)
The relationship between the Fourier coefficients from (8) and the amplitudes and phase shifts of this equivalent form (13) is presented below. Stating this relationship for a general dimension $m$ requires a more descriptive framework than the sequential indexing (e.g., $a_{1}^{(n)},a_{2}^{(n)}$) suitable for the two–dimensional case. We aim to calculate the coefficient $a_{k}^{(\boldsymbol{n})}$ corresponding to the basis function
@@
where the set $\mathcal{I}_{k}\subseteq\{1,\dots,m\}$ and its complement $\mathcal{I}_{k}^{c}=\{1,\dots,m\}\setminus\mathcal{I}_{k}$ specify the indices of the sinesine and cosine terms, respectively, for the basis corresponding to the index $k\in\{1,\dots,2^{m}\}$. The index set $\mathcal{I}_{k}$ is determined from the sequential index $k$ using the mm–bit binary representation of k1k-1: an index $j$ is included in $\mathcal{I}_{k}$ if and only if the $j^{th}$ bit from the left is $1$. Given the amplitudes $A_{i}^{(\boldsymbol{n})}$ and the phase shifts $\phi_{i}^{(\boldsymbol{n})}$, the separable product–basis Fourier coefficients $a_{k}^{(\boldsymbol{n})}$, are related by
@@ 4 Model Construction and Architecture
Definition 2 (Single-Layer Feedforward Neural Network).
A single–hidden–layer NN with an input vector $\boldsymbol{x}\in\mathbb{R}^{m}$, $L$ hidden neurons, and a single output $\widehat{f}(\boldsymbol{x})\in\mathbb{R}$ is a function defined as:
$\widehat{f}(\boldsymbol{x})=g\left(\sum_{j=1}^{L}w_{j}^{\text{out}}\cdot\sigma(\boldsymbol{w}_{j}^{\text{in}}\cdot\boldsymbol{x}-b_{j})\right)$
where $\boldsymbol{w}_{j}^{\text{in}}$ is the input–to–hidden weight vector, $b_{j}$ is the hidden neuron bias, $\sigma(\cdot)$ is the hidden layer activation function, $w_{j}^{\text{out}}$ is the hidden–to–output weight, and $g(\cdot)$ is the output layer activation function.
Definition 3 (Fourier Learning Machine (FLM)).
The FLM is composed of $N$ parallel sub–networks, each consisting of a complete single–layer NN conforming to Definition 2, with its hidden activation function $\sigma:x\mapsto\cos(x)$, and its output activation, $g:x\mapsto x$. The weight applied to the input to the top neuron in the hidden–layer of each sub–network is set to be a row vector $\boldsymbol{n}$. To each subsequent neuron in the hidden-layer, the same weight vector is used, with the respective sign of its components determined by the mm–Lexi Sign Matrix. The final output of the entire FLM is the ordinary sum of the outputs from all these individual sub–networks.
Next, we present two instances of possible sub–networks, first for the two-input case and then for the general $m$-input case.
Figure 2: FLM with two inputs. The sub–network associated with frequency pair $\boldsymbol{n}$ (Left) and the main network with NN sub–networks (Right).
Example 4.1 (FLM sub–networks with Two Inputs).
An FLM with two inputs is composed of $N$ sub–networks, each parameterized by a learnable weight vector $\boldsymbol{n}$, with input $(x_{1},x_{2})$, two hidden–neurons with a cosinecosine activation function and bias. The vector $\boldsymbol{n}=(n_{1},n_{2})$ contains the weights from the input–layer to the first hidden–neuron. For the second hidden–neuron, the same weights are reused except that the second element $n_{2}$ has opposite sign. The output of the sub–network is the weighted sum of outputs of the two cosinecosine–activated hidden–neurons, i.e.,
$\widehat{H}_{\boldsymbol{n}}$ $=\widehat{A}_{1}^{(\boldsymbol{n})}\cos(n_{1}x_{1}+n_{2}x_{2}-b_{1}^{(\boldsymbol{n})})+\widehat{A}_{2}^{(\boldsymbol{n})}\cos(n_{1}x_{1}-n_{2}x_{2}-b_{2}^{(\boldsymbol{n})}).$
@@
$\widehat{f}(x_{1},x_{2})=\sum_{\boldsymbol{n}\in\mathcal{N}}\widehat{H}_{\boldsymbol{n}}(x_{1},x_{2}),$ (14)
where $\mathcal{N}$ is the set of learnable frequency vectors, one from each sub–network. (14) has the same form of the Fourier series in (10). The hidden–to–output layer weights $\widehat{A}_{i}^{(\boldsymbol{n})}$ of the sub–network can be interpreted as the amplitudes $A_{i}^{(\boldsymbol{n})}$and the biases $b_{i}^{(\boldsymbol{n})}$ as phase shifts $\phi_{i}^{(\boldsymbol{n})}$, respectively, for $i=1,2$. Fig. 2 illustrates a FLM sub–network with two inputs and the FLM itself.
Now we discuss the structure of the FLM sub–networks in general for mm–inputs and explain the interpretation of the multidimensional FLM.
Example 4.2 (FLM Sub–Networks with $m$ inputs).
An mm–input FLM has $N$ parallel sub–networks, each of which is associated with a weight vector $\boldsymbol{n}=(n_{1},n_{2},\dots,n_{m})$. It has inputs $x_{1},x_{2},\dots,x_{m}$ and $l=2^{m-1}$ hidden nodes with cosinecosine activation function and biases $b_{i}^{(\boldsymbol{n})}$ for $i=1,\dots,l$. The input–to–hidden layer weights for the $i^{th}$ node of the sub–network is the Hadamard (element wise) product $\boldsymbol{e}^{(i)}\odot\boldsymbol{n}$ where the row vector $\boldsymbol{e}^{(i)}$ is the $i^{th}$ row of the mm–Lexi Sign Matrix $S$ (12). An illustration of the mm–dimensional FLM sub–network is presented in Fig. 3.
@@
$\widehat{H}_{\boldsymbol{n}}$ $=\sum_{i=1}^{l}\widehat{A}_{i}^{(\boldsymbol{n})}\cos\left[\left(e_{1}^{(i)}n_{1}x_{1}+e_{2}^{(i)}n_{2}x_{2}+\dots+e_{m}^{(i)}n_{m}x_{m}\right)-b_{i}^{(\boldsymbol{n})}\right]$
$=\sum_{i=1}^{l}\widehat{A}_{i}^{(\boldsymbol{n})}\cos\left((\boldsymbol{e}^{(i)}\odot\boldsymbol{n})\cdot\boldsymbol{x}-b_{i}^{(\boldsymbol{n})}\right).$
@@
$\widehat{f}(x_{1},x_{2},\dots,x_{m})=\sum_{\boldsymbol{n}\in\mathcal{N}}\widehat{H}_{\boldsymbol{n}}(x_{1},x_{2},\dots,x_{m}),$
where $\mathcal{N}$ is the finite set containing the learnable frequency vector from each sub–network. Notice that $\widehat{H}_{\boldsymbol{n}}$ matches the form of $H_{\boldsymbol{n}}$ and, consequently, $\widehat{f}$ with the Fourier series $\widehat{f}_{NH}$ itself in (13). The hidden–to–output layer weights $\widehat{A}_{i}^{(\boldsymbol{n})}$ of the sub–network can be interpreted as the amplitudes $A_{i}^{(\boldsymbol{n})}$ and the biases $b_{i}^{(\boldsymbol{n})}$ as phase shifts $\phi_{i}^{(\boldsymbol{n})}$, respectively for $i=1,\dots,l$. Thus each sub–network with $2^{m-1}$ neurons explicitly generates a composition of $2^{m-1}$ mm–dimensional phase–shifted cosinescosines or implicitly $2^{m}$ separable Fourier basis with a specific frequency vector. The relative contributions of each individual Fourier component are learned through the sub–network biases and output layer weights. Although the network does not perform explicit projection onto the orthogonal bases, it learns both the amplitude and phase of each learned basis function, enabling a structured and interpretable approximation of the target map.
We have demonstrated in this section that a MLP–like, feedforward NN with traditional node operations can represent a mm–dimensional nonharmonic Fourier series. In the next section, we compare the performance of the our FLM model against other NN models.
@@ 5 Computational Results
In this section we test and compare the performance of our proposed FLM model in the computation of numerical solutions of three PDEs and an OCP arising from an odd-circulant evolutionary game. Additionally, we design a series of computational experiments aimed at comparing the performance of our proposed model against other traditionally employed NN architectures.
The PDEs considered herein are: the one–dimensional heat equation, the two–dimensional Poisson equation, and the generalized Black–Scholes model. Besides the emerging interest in machine learning techniques for the solution of PDEs, demonstrated, for example, in [18, 21], the solution of OCPs by such methods was also recently brought to attention (see, e.g., [22, 23, 24, 25, 26]). Thus, we consider in this paper an OCP arising from the evolutionary dynamics of a Rock–Paper–Scissors game, as studied in [27]. Its nonlinear dynamics alongside its state space configuration make this OCP particularly challenging to be solved analytically, so our solution shall be tested against an (indirect) numerical solution based on the Pontryagin’s Minimum Principle (PMP). While traditional techniques like Finite Elements and Finite Differences are well–established methods, the recent rise in interest on machine learning-based techniques has introduced a new class of powerful, mesh–free solvers based on NNs [28]. In this section, we demonstrate the efficacy of our proposed FLMs in solving several benchmark PDEs, comparing its performance against other prominent NN architectures and a class of OCPs.
@@ 5.1 Numerical Experiments for Partial Differential Equations
$\mathcal{F}[u](\boldsymbol{x},t)$ $=0,$ $(\boldsymbol{x},t)\in\Omega\times(0,T],$
$u(\boldsymbol{x},0)\quad$ $=u_{0}(\boldsymbol{x}),$ $\boldsymbol{x}\in\Omega,$
$\mathcal{B}[u](\boldsymbol{x},t)$ $=g(\boldsymbol{x},t),$ $(\boldsymbol{x},t)\in\partial\Omega\times(0,T].$
Here, $\partial\Omega$ is the spatial boundary of $\Omega$, $\mathcal{F}[\cdot]$ and $\mathcal{B}[\cdot]$ denote general (possibly nonlinear) differential operators, $u(\boldsymbol{x},t)$ is the unknown solution, and $u_{0}(\boldsymbol{x})$ and $g(\boldsymbol{x},t)$ define the initial and boundary conditions, respectively.
We approximate the solution $u(\boldsymbol{x},t)$ using a FLM $u_{\Theta}(\boldsymbol{x},t)$ parameterized by its weights and biases, collectively denoted as $\Theta$. The network is trained by minimizing a composite loss function $\mathcal{L}(\Theta)$ (such as in [29]), that penalizes deviations from the governing PDE and its associated initial and boundary conditions
@@
$\mathcal{L}_{\text{IC}}$ $=\frac{1}{N_{\text{IC}}}\sum_{i=1}^{N_{\text{IC}}}\big{\lvert}u_{\Theta}(\boldsymbol{x}_{i},0)-u_{0}(\boldsymbol{x}_{i})\big{\rvert}^{2},$
$\mathcal{L}_{\text{BC}}$ $=\frac{1}{N_{\text{BC}}}\sum_{j=1}^{N_{\text{BC}}}\big{\lvert}\mathcal{B}[u_{\Theta}](\boldsymbol{x}_{j},t_{j})-g(\boldsymbol{x}_{j},t_{j})\big{\rvert}^{2},$
$\mathcal{L}_{\text{PDE}}$ $=\frac{1}{N_{\text{PDE}}}\sum_{k=1}^{N_{\text{PDE}}}\big{\lvert}\mathcal{F}[u_{\Theta}](\boldsymbol{x}_{k},t_{k})\big{\rvert}^{2}.$
Each component of the loss function (15) is a mean square error (MSE) computed over a distinct set of collocation points, defining the training set. These points are randomly sampled from their corresponding domains: $\{\boldsymbol{x}_{i}\}$ from the spatial domain $\Omega$ at $t=0$, $\{(\boldsymbol{x}_{j},t_{j})\}$ from the boundary $\partial\Omega\times[0,T]$, and $\{(\boldsymbol{x}_{k},t_{k})\}$ from the interior $\Omega\times(0,T]$. The integers $N_{\text{IC}}$, $N_{\text{BC}}$, and $N_{\text{PDE}}$ are the number of the training points used to enforce the initial, boundary, and PDE constraints in the loss function, respectively. The differential operators $\mathcal{F}[\cdot]$ and $\mathcal{B}[\cdot]$ are evaluated on these points using automatic differentiation, and the loss function is then minimized with adaptive momentum gradient descent algorithm (ADAM) [30].
@@
SIREN–1 SIREN with 1 hidden layer (only the SIREN layer with $\omega_{0}=1$)
SIREN–2 SIREN with 2 hidden layers (the SIREN layer with $\omega_{0}=1$ and one hidden layer)
SIREN–3 SIREN with 3 hidden layers (the SIREN layer with $\omega_{0}=1$ and tow hidden layers)

### v3

@@ 3.2 Cosine Phase-Shifted Basis for $m$-dimensional Input
In a similar fashion as in the two–dimensional case, $H_{\bm{n}}$ represents the linear combination of $l=2^{m-1}$ phase–shifted cosinescosines, which is equivalent to a linear combination of $2^{m}$ separable Fourier basis in Equation 8 and can be written as
$H_{\bm{n}}=\sum_{i=1}^{l}A_{i}^{(\bm{n})}\cos\left((\bm{e}^{(i)}\odot\bm{n})\cdot\bm{x}-\phi_{i}^{(\bm{n})}\right),$
where $\bm{e}^{(i)}$ is the $i^{th}$ row of the $m$-Lexi Sign Matrix $S$ and $\bm{e}^{(i)}\odot\bm{n}$ represents its Hadamard (element wise) product with the frequency vector $\bm{n}$. The $m$-dimensional Fourier series can be expressed in a single compact summation as
$\widehat{f}_{NH}(\bm{x})=\sum_{\bm{n}\in\mathcal{N}}H_{\bm{n}}.$ (13)
The relationship between the Fourier coefficients from Equation 8 and the amplitudes and phase shifts of this equivalent form Equation 13 is presented below. Stating this relationship for a general dimension $m$ requires a more descriptive framework than the sequential indexing (e.g., $a_{1}^{(n)},a_{2}^{(n)}$) suitable for the two–dimensional case. We aim to calculate the coefficient $a_{k}^{(\bm{n})}$ corresponding to the basis function
@@
where the set $\mathcal{I}_{k}\subseteq\{1,\dots,m\}$ and its complement $\mathcal{I}_{k}^{c}=\{1,\dots,m\}\setminus\mathcal{I}_{k}$ specify the indices of the sinesine and cosine terms, respectively, for the basis corresponding to the index $k\in\{1,\dots,2^{m}\}$. The index set $\mathcal{I}_{k}$ is determined from the sequential index $k$ using the mm–bit binary representation of k1k-1: an index $j$ is included in $\mathcal{I}_{k}$ if and only if the $j^{th}$ bit from the left is $1$. Given the amplitudes $A_{i}^{(\bm{n})}$ and the phase shifts $\phi_{i}^{(\bm{n})}$, the separable product–basis Fourier coefficients $a_{k}^{(\bm{n})}$, are related by
@@ 4 Model Construction and Architecture
Definition 2 (Single-Layer Feedforward Neural Network)
A single–hidden–layer NN with an input vector $\bm{x}\in\mathbb{R}^{m}$, $l$ hidden neurons, and a single output $\widehat{f}(\bm{x})\in\mathbb{R}$ is a function defined as:
$\widehat{f}(\bm{x})=g\left(\sum_{j=1}^{l}w_{j}^{\text{out}}\cdot\sigma(\bm{w}_{j}^{\text{in}}\cdot\bm{x}-b_{j})\right)$
where $\bm{w}_{j}^{\text{in}}$ is the input–to–hidden weight vector, $b_{j}$ is the hidden neuron bias, $\sigma(\cdot)$ is the hidden layer activation function, $w_{j}^{\text{out}}$ is the hidden–to–output weight, and $g(\cdot)$ is the output layer activation function.
Definition 3 (Fourier Learning Machine (FLM))
The FLM is composed of $N$ parallel sub–networks, each consisting of a complete single–layer NN conforming to Definition 2, with its hidden activation function $\sigma:x\mapsto\cos(x)$, and its output activation, $g:x\mapsto x$. The weight applied to the input to the top neuron in the hidden–layer of each sub–network is set to be a row vector $\bm{n}$. To each subsequent neuron in the hidden-layer, the same weight vector is used, with the respective sign of its components determined by the mm–Lexi Sign Matrix. The final output of the entire FLM is the ordinary sum of the outputs from all these individual sub–networks.
Next, we present three instances of possible sub–networks, first for the two–input and three–input cases and then for the general $m$-input case.
Example 4.1 (FLM Sub–networks with Two Inputs)
An FLM with two inputs is composed of $N$ sub–networks, each parameterized by a learnable weight vector $\bm{n}$, with input $(x_{1},x_{2})$, two hidden–neurons with a cosinecosine activation function and bias. The vector $\bm{n}=(n_{1},n_{2})$ contains the weights from the input–layer to the first hidden–neuron. For the second hidden–neuron, the same weights are reused except that the second element $n_{2}$ has opposite sign. The output of the sub–network is the weighted sum of outputs of the two cosinecosine–activated hidden–neurons, i.e.,
$\widehat{H}_{\bm{n}}$ $=\widehat{A}_{1}^{(\bm{n})}\cos(n_{1}x_{1}+n_{2}x_{2}-b_{1}^{(\bm{n})})+\widehat{A}_{2}^{(\bm{n})}\cos(n_{1}x_{1}-n_{2}x_{2}-b_{2}^{(\bm{n})}).$
@@
$\widehat{f}(x_{1},x_{2})=\sum_{\bm{n}\in\mathcal{N}}\widehat{H}_{\bm{n}}(x_{1},x_{2}),$ (14)
where $\mathcal{N}$ is the set of learnable frequency vectors, one from each sub–network. Equation 14 has the same form of the Fourier series in Equation 10. The hidden–to–output layer weights $\widehat{A}_{i}^{(\bm{n})}$ of the sub–network can be interpreted as the amplitudes $A_{i}^{(\bm{n})}$and the biases $b_{i}^{(\bm{n})}$ as phase shifts $\phi_{i}^{(\bm{n})}$, respectively, for $i=1,2$. Fig. 2 illustrates a FLM sub–network with two inputs and the FLM itself.
Figure 2: FLM with two inputs. The sub–network associated with frequency pair $\bm{n}$ (Left) and the main network with NN sub–networks (Right).
Additionally, a FLM with 3 inputs is shown in Fig. 3 alongside its underlying $S^{(3)}$ matrix. Now we discuss the structure of the FLM sub–networks in general for mm–inputs and explain the interpretation of the multidimensional FLM.
Figure 3: $(a)$ FLM with 3 inputs and the $S^{(3)}$ matrix defining the sign of the frequency components $n_{1}$, $n_{2}$ and $n_{3}$. These components are illustrated, respectively, in (b)(d)(b)-(d).
Example 4.2 (FLM Sub–Networks with $m$ inputs)
An mm–input FLM has $N$ parallel sub–networks, each of which is associated with a weight vector $\bm{n}=(n_{1},n_{2},\dots,n_{m})$. It has inputs $x_{1},x_{2},\dots,x_{m}$ and $l=2^{m-1}$ hidden nodes with cosinecosine activation function and biases $b_{i}^{(\bm{n})}$ for $i=1,\dots,l$. The input–to–hidden layer weights for the $i^{th}$ node of the sub–network is the Hadamard (element wise) product $\bm{e}^{(i)}\odot\bm{n}$ where the row vector $\bm{e}^{(i)}$ is the $i^{th}$ row of the mm–Lexi Sign Matrix $S^{(m)}$ (Equation 12). An illustration of the mm–dimensional FLM sub–network is presented in Fig. 4.
@@
$\widehat{H}_{\bm{n}}$ $=\sum_{i=1}^{l}\widehat{A}_{i}^{(\bm{n})}\cos\left[\left(e_{1}^{(i)}n_{1}x_{1}+e_{2}^{(i)}n_{2}x_{2}+\dots+e_{m}^{(i)}n_{m}x_{m}\right)-b_{i}^{(\bm{n})}\right]$
$=\sum_{i=1}^{l}\widehat{A}_{i}^{(\bm{n})}\cos\left((\bm{e}^{(i)}\odot\bm{n})\cdot\bm{x}-b_{i}^{(\bm{n})}\right).$
@@
$\widehat{f}(x_{1},x_{2},\dots,x_{m})=\sum_{\bm{n}\in\mathcal{N}}\widehat{H}_{\bm{n}}(x_{1},x_{2},\dots,x_{m}),$
where $\mathcal{N}$ is the finite set containing the learnable frequency vector from each sub–network. Notice that $\widehat{H}_{\bm{n}}$ matches the form of $H_{\bm{n}}$ and, consequently, $\widehat{f}$ with the Fourier series $\widehat{f}_{NH}$ itself in Equation 13. The hidden–to–output layer weights $\widehat{A}_{i}^{(\bm{n})}$ of the sub–network can be interpreted as the amplitudes $A_{i}^{(\bm{n})}$ and the biases $b_{i}^{(\bm{n})}$ as phase shifts $\phi_{i}^{(\bm{n})}$, respectively for $i=1,\dots,l$. We note that the $2^{m-1}$ neurons per sub–network grow exponentially with $m$, which is an inherent consequence of representing an mm–dimensional Fourier basis in separable form. Thus, each sub–network with $2^{m-1}$ neurons explicitly generates a composition of $2^{m-1}$ mm–dimensional phase–shifted cosinescosines or implicitly $2^{m}$ separable Fourier basis with a specific frequency vector. The relative contributions of each individual Fourier component are learned through the sub–network biases and output layer weights. Although the network does not perform explicit projection onto the orthogonal bases, it learns both the amplitude and phase of each learned basis function, enabling a structured and interpretable approximation of the target map.
We have demonstrated in this section that a MLP–like, feedforward NN with traditional node operations can represent a mm–dimensional nonharmonic Fourier series. In the next section, we compare the performance of our FLM model against other NN models.
@@ 5 Computational Results
In this section we test and compare the performance of our proposed FLM model in the computation of numerical solutions of four PDEs and an OCP arising from an odd–circulant evolutionary game. Additionally, we design a series of computational experiments aimed at comparing the performance of our proposed model against other traditionally employed NN architectures.
The PDEs considered herein are: the one–dimensional heat equation, the two–dimensional Poisson equation, the generalized Black–Scholes model, and the one–dimensional inviscid Burgers’ equation. Besides the emerging interest in machine learning techniques for the solution of PDEs, demonstrated, for example, in Ngom and Marin (2021); Ma et al. (2022), the solution of OCPs by such methods was also recently brought to attention (see, e.g., Nicolosi et al. (2022); Nicolosi and Griffin (2023); Nicolosi et al. (2025); Rubel and Nicolosi (2025); Lundqvist and Oliveira (2025)). Thus, we consider in this paper an OCP arising from the evolutionary dynamics of a Rock–Paper–Scissors game, as studied in Griffin and Fan (2022). Its nonlinear dynamics alongside its state space configuration make this OCP particularly challenging to be solved analytically, so our solution shall be tested against an (indirect) numerical solution based on the Pontryagin’s Minimum Principle (PMP). While traditional techniques like Finite Elements and Finite Differences are well–established methods, the recent rise in interest on machine learning-based techniques has introduced a new class of powerful, mesh–free solvers based on NNs (Raissi et al., 2019). In this section, we demonstrate the efficacy of our proposed FLMs in solving several benchmark PDEs and a class of OCPs.
@@ 5.1 Numerical Experiments for Partial Differential Equations
$\mathcal{F}[u](\bm{x},t)$ $=0,$ $(\bm{x},t)\in\Omega\times(0,T],$
$u(\bm{x},0)\quad$ $=u_{0}(\bm{x}),$ $\bm{x}\in\Omega,$
$\mathcal{B}[u](\bm{x},t)$ $=g(\bm{x},t),$ $(\bm{x},t)\in\partial\Omega\times(0,T].$
Here, $\partial\Omega$ is the spatial boundary of $\Omega$, $\mathcal{F}[\cdot]$ and $\mathcal{B}[\cdot]$ denote general (possibly nonlinear) differential operators, $u(\bm{x},t)$ is the unknown solution, and $u_{0}(\bm{x})$ and $g(\bm{x},t)$ define the initial and boundary conditions, respectively.
We approximate the solution $u(\bm{x},t)$ using a FLM $u_{\Theta}(\bm{x},t)$ parameterized by its weights and biases, collectively denoted as $\Theta$. The network is trained by minimizing a composite loss function $\mathcal{L}(\Theta)$ (such as in Cuomo et al. (2022)), that penalizes deviations from the governing PDE and its associated initial and boundary conditions
@@
$\mathcal{L}_{\text{IC}}$ $=\frac{1}{N_{\text{IC}}}\sum_{i=1}^{N_{\text{IC}}}\big\lvert u_{\Theta}(\bm{x}_{i},0)-u_{0}(\bm{x}_{i})\big\rvert^{2},$
$\mathcal{L}_{\text{BC}}$ $=\frac{1}{N_{\text{BC}}}\sum_{j=1}^{N_{\text{BC}}}\big\lvert\mathcal{B}[u_{\Theta}](\bm{x}_{j},t_{j})-g(\bm{x}_{j},t_{j})\big\rvert^{2},$
$\mathcal{L}_{\text{PDE}}$ $=\frac{1}{N_{\text{PDE}}}\sum_{k=1}^{N_{\text{PDE}}}\big\lvert\mathcal{F}[u_{\Theta}](\bm{x}_{k},t_{k})\big\rvert^{2}.$
Each component of the loss function Equation 15 is a mean square error (MSE) computed over a distinct set of collocation points, defining the training set. These points are randomly sampled from their corresponding domains: $\{\bm{x}_{i}\}$ from the spatial domain $\Omega$ at $t=0$, $\{(\bm{x}_{j},t_{j})\}$ from the boundary $\partial\Omega\times[0,T]$, and $\{(\bm{x}_{k},t_{k})\}$ from the interior $\Omega\times(0,T]$. The integers $N_{\text{IC}}$, $N_{\text{BC}}$, and $N_{\text{PDE}}$ are the number of the training points used to enforce the initial, boundary, and PDE constraints in the loss function, respectively. The differential operators $\mathcal{F}[\cdot]$ and $\mathcal{B}[\cdot]$ are evaluated on these points using automatic differentiation, and the loss function is then minimized with adaptive momentum gradient descent algorithm (ADAM) (Kingma and Ba, 2014).
@@
SIREN–1 SIREN with 1 hidden layer ($\omega_{0}=1$)
SIREN–2 SIREN with 2 hidden layers ($\omega_{0}=1$)
SIREN–3 SIREN with 3 hidden layers ($\omega_{0}=1$)